FINITE GROUPS ACTING LINEARLY: HOCHSCHILD COHOMOLOGY AND THE CUP PRODUCT

ANNE V. SHEPLER AND SARAH WITHERSPOON

Abstract. When a finite group acts linearly on a complex vector space, the natural semi-direct product of the group and the polynomial ring over the space forms a skew group algebra. This algebra plays the role of the coordinate ring of the resulting orbifold and serves as a substitute for the ring of invariant polynomials from the viewpoint of geometry and physics. Its Hochschild cohomology predicts various Hecke algebras and deformations of the orbifold. In this article, we investigate the ring structure of the Hochschild cohomology of the skew group algebra. We show that the cup product coincides with a natural smash product, transferring the cohomology of a group action into a group action on cohomology. We express the algebraic structure of Hochschild cohomology in terms of a partial order on the group (modulo the kernel of the action). This partial order arises after assigning to each group element the codimension of its fixed point space. We describe the algebraic structure for Coxeter groups, where this partial order is given by the reflection length function; a similar combinatorial description holds for an infinite family of complex reflection groups.

1. Introduction

Physicists often regard space as a Calabi-Yau manifold M endowed with symmetries forming a group G. The orbifold M/G is not smooth in general, and they regularly shift focus from the orbifold to a desingularisation and its coordinate ring. In the affine case, we take M to be a finite dimensional, complex vector space V upon which a finite group G acts linearly. The orbifold V/G may then be realized as an algebraic variety whose coordinate ring is the ring of invariant polynomials S(V^{∗})^{G} on the dual space V^{∗} (see Harris [18]). The variety V/G is nonsingular exactly when the action of G on V is generated by reflections. When the orbifold V/G is singular, we seek to replace the space of invariant polynomials with a natural algebra attached to V/G playing the role of a coordinate ring.

Date: July 21, 2009. The first author was partially supported by NSF grants #DMS-0402819 and #DMS-0800951

and a research fellowship from the Alexander von Humboldt Foundation. The second author was partially supported by NSA grant #H98230-07-1-0038 and NSF grant #DMS-0800832. Both authors were jointly supported by Advanced Research Program Grant 010366-0046-2007 from the Texas Higher Education Coordinating Board.

In situations arising naturally in physics, a noncommutative substitute for the commutative invariant ring S(V^{∗})^{G} is provided by the skew group algebra

S(V^{∗})#G = S(V^{∗}) ⋊ G,

the natural semi-direct product of G with the symmetric algebra S(V^{∗}). One resolves the singularities of V/G with a smooth Calabi-Yau variety X whose coordinate ring behaves as the skew group algebra. Indeed, the McKay equivalence implies (in certain settings) that Hochschild cohomology sees no difference:

HH^{}(S(V^{∗})#G) ≅ HH^{}(O_{X})

as algebras under both the cup product and Lie bracket. Note that the Hochschild cohomology HH^{}(S(V^{∗})#G) can also be used to recover the orbifold (or stringy) cohomology H_{orb}(V/G) (which is isomorphic to the singular cohomology of X). (See [3], [6], [7], [15], [19].) The cohomology and deformations of S(V)#G appear in various other areas of mathematics as well—for example, combinatorics, representation theory, Lie theory, noncommutative algebra, and invariant theory (see, for example, Etingof and Ginzburg [11]).

In this paper, we consider any finite group G acting linearly on V and explore the rich algebraic structure of the Hochschild cohomology of S(V)#G under the cup product. This structure is interesting not only in its own right, but also because of possible applications in algebra and representation theory. For example, the graded Lie bracket, which predicts potential deformations (like sympletic reflection algebras and graded Hecke algebras), is a graded derivation on Hochschild cohomology with respect to the cup product. The representation theory of finite dimensional algebras provides an application of the cup product in a different setting: Often, one may associate an algebraic variety to each module over the algebra using the ring structure of its Hochschild cohomology; the collection of such varieties provide a coarse invariant of the representation theory of the algebra (see, for example, Snashall and Solberg [26]).

For any algebra A over a field k, Hochschild cohomology HH^{}(A) is the space Ext_{A⊗A^{op}}(A, A), which is a Gerstenhaber algebra under the two compatible operations, cup product and bracket. Both operations are defined initially on the bar resolution, a natural A ⊗ A^{op}-free resolution of A. For A = S(V)#G, an explicit description of Hochschild cohomology HH^{}(A) arises not from the bar resolution of A, but instead from a Koszul resolution: HH^{}(S(V)#G) is isomorphic to the G-invariant subalgebra of HH^{}(S(V), S(V)#G), which is computed with a Koszul resolution of S(V). We analyze the cup product on HH^{}(S(V)#G) by taking advantage of these two different manifestations of cohomology arising from two different resolutions. We show that the cup product on HH^{}(S(V)#G) may be written as the cup product on S(V) twisted by the action of the group. This perspective yields convenient descriptions of the ring structure of cohomology. (We study the Gerstenhaber bracket in a future work.)

Over the real numbers, the cup product on HH^{}(S(V)#G) has been studied in related settings (e.g., see [22]). But note that this analysis of Hochschild cohomology does not readily extend to the complex numbers: Over the real numbers, V and V^{∗} are naturally G-isomorphic, which simplifies several aspects of the theory. We develop the theory in the richer setting of complex affine space. Note that our results actually hold over any field containing the eigenvalues of the action of G on V in which |G| is invertible. Anno [1] also gave a cup product formula in the geometric setting over fairly general fields; we give a natural interpretation of the resulting ring structure from a purely algebraic and combinatorial point of view. In this paper, we do not assume that G acts symplectically or even faithfully, as there are interesting applications in which it does not (see [9, 10, 24]).

In Section 2, we define the “codimension poset” which arises from assigning to each group element the codimension of its fixed point space. We also posit a few observations needed later on the geometry of finite group actions. In Sections 3 and 4, we establish definitions and notation and recall necessary facts about the bar and Koszul resolutions. We review the structure of HH^{}(S(V)#G) as a graded vector space as well, originally found independently by Farinati [13] and Ginzburg and Kaledin [15] for faithful group actions. In Section 5, we define a combinatorial map in terms of Demazure operators and quantum differentiation which allows for conversion between complexes. This combinatorial conversion map (introduced in [25]) induces isomorphisms on cohomology.

In Sections 6 and 7, we transform the “cohomology of a group action” into a “group action on cohomology”. One may first take Hochschild cohomology and then form the skew group algebra, or one may reverse this order of operation. In Section 6, we compare resulting algebras for S(V):

HH^{}(S(V)#G) versus HH^{}(S(V))#G versus HH^{}(S(V)) ⊗ ℂ[G] .

Since HH^{}(S(V)#G) is the G-invariant subalgebra of HH^{}(S(V), S(V)#G), we focus on this latter ring. We show that the smash product on HH^{}(S(V))#G induces a smash product on HH^{}(S(V), S(V)#G), which we then view as an algebra under three operations:

• the cup product ⌣ induced from the bar resolution of S(V),

• the smash product ⋄ induced from HH^{}(S(V))#G, and

• the usual multiplication in the tensor algebra product HH^{}(S(V)) ⊗ ℂ[G].

In Section 7 (see Theorem 7.1), we show that these three algebraic operations coincide. This yields a simple formula for the cup product on HH^{}(S(V), S(V)#G) (see Theorem 7.3; cf. Anno [1]) and implies that HH^{}(S(V)#G) is isomorphic to an algebra subquotient of HH^{}(S(V))#G (see Corollary 7.6). These results express the cup product on HH^{}(S(V)#G) (at the cochain level) as the cup product on HH^{}(S(V)) twisted by the group G.

In Section 8, we identify an interesting graded subalgebra (the “volume subalgebra”) of HH^{}(S(V), S(V)#G) whose dimension is the order of G. When G is a subgroup of the symplectic group Sp(V), its G-invariant subalgebra is isomorphic to the cohomology of the G-invariant subalgebra of the Weyl algebra. In this case, it is also isomorphic to the orbifold cohomology of V/G. (See Remark 8.2.) We thus display the orbifold cohomology H_{orb}(V/G) as a natural subalgebra of HH^{}(S(V)#G).

In Section 9, we describe generators of cohomology (as an algebra) via the codimension poset. The partial order is defined on G modulo the kernel of its action on V. We view division in the volume algebra as a purely geometric construction by interpreting results in terms of this poset. Generators for the Hochschild cohomology HH^{}(S(V), S(V)#G) arise from minimal elements in the poset (with the identity removed); see Corollaries 9.3 and 9.4.

In Section 10, we explore implications for reflection groups. The reflection length of a group element is the length of a shortest word expressing that element as a product of reflections. For Coxeter groups and many complex reflection groups, reflection length coincides with the codimension of the fixed point space. In this case, the codimension poset appears as a well-studied poset (arising from reflection length) in the theory of reflection groups. We show (in Corollary 10.6) that for Coxeter groups and many other complex reflection groups, the Hochschild cohomology HH^{}(S(V), S(V)#G) is generated as an algebra in homological degrees 0 and 1, in analogy with the Hochschild-Kostant-Rosenberg Theorem for smooth commutative algebras (such as S(V)).

Finally, in Section 11, we return to HH^{}(S(V)#G) as the G-invariant subalgebra of HH^{}(S(V), S(V)#G). We point out the simple cup product structure in cohomological degrees 0 and 1. In Theorem 11.4, we use standard group-theoretic techniques (Green/Mackey functors and transfer maps) to describe the product on HH^{}(S(V)#G). We use the natural symplectic action of the symmetric group to give a nontrivial example.

2. Poset and volume forms

We begin by collecting several geometric observations needed later. Let G be a finite group and V a ℂG-module of finite dimension n. Denote the image of v ∈ V under the action of g ∈ G by ^{g}v. We work with the induced group action on maps: For any function θ and any element h ∈ GL(V) acting on its domain and range, we define the map ^{h}θ by (^{h}θ)(v) := ^{h}(θ(^{h^{−1}}v)). Let V^{∗} denote the contragredient (or dual) representation. For any basis v_{1}, …, v_{n} of V, let v_{1}^{∗}, …, v_{n}^{∗} be the dual basis of V^{∗}. Let V^{G} be the set of G-invariants in V: V^{G} = {v ∈ V : ^{g}v = v for all g ∈ G}. For any g ∈ G, set V^{g} = {v ∈ V : ^{g}v = v}, the fixed point set of g in V. Since G is finite, we assume G acts by isometries on V (i.e., G preserves a given inner product). All tensor and exterior products will be taken over ℂ unless otherwise indicated.

We regard G modulo the kernel of its action on V as a poset using the following lemma, which is no surprise (see [4] or [15], for example). Note that (V^{g})^{⊥} = im(1 − g) for all g in G.

Lemma 2.1. Let g, h ∈ G. The following are equivalent:

(i) (V^{g})^{⊥} ∩ (V^{h})^{⊥} = 0,

(ii) V^{g} + V^{h} = V,

(iii) codim V^{g} + codim V^{h} = codim V^{gh},

(iv) (V^{g})^{⊥} ⊕ (V^{h})^{⊥} = (V^{gh})^{⊥}. Any of these properties implies that V^{g} ∩ V^{h} = V^{gh}.

Proof. Taking orthogonal complements yields the equivalence of (i) and (ii).

Assume (ii) holds and write u ∈ V^{gh} as v + w where v ∈ V^{g} and w ∈ V^{h}. Then ^{h}v +^{h}w = ^{g^{−1}}v +^{g^{−1}}w and ^{h}v + w = v +^{g^{−1}}w, i.e., ^{(h−1)}v = ^{(g^{−1}−1)}w. But (V^{h})^{⊥} = im(h − 1) while (V^{g})^{⊥} = (V^{g^{−1}})^{⊥} = im(g^{−1} − 1), and the intersection of these spaces is 0 by (i), hence ^{(h−1)}v = 0 = ^{(g^{−1}−1)}w. Therefore, v ∈ V^{h}, w ∈ V^{g}, and u ∈ V^{g} ∩ V^{h}, and thus V^{gh} ⊂ V^{g} ∩ V^{h}. The reverse inclusion is immediate, hence V^{gh} = V^{g} ∩ V^{h}. We take orthogonal complements and observe that (V^{gh})^{⊥} = (V^{g})^{⊥} ⊕ (V^{h})^{⊥}. A dimension count then gives (iii).

To show (iii) implies (iv), note (V^{gh})^{⊥} ⊂ (V^{g})^{⊥} + (V^{h})^{⊥} since V^{g} ∩ V^{h} ⊂ V^{gh}. By (iii), this containment is forced to be an equality and the sum is direct: (V^{g})^{⊥} ⊕ (V^{h})^{⊥} = (V^{gh})^{⊥}. As (iv) trivially implies (i), we are finished. □

Let K be the kernel of the representation of G acting on V:

K := {k ∈ G : V^{k} = V }.

Definition 2.2. Define a binary relation ≤ on G by g ≤ h whenever

codim(V^{g}) + codim(V^{g^{−1}h}) = codim(V^{h}).

By Lemma 2.1, this codimension condition holds exactly when

(V^{g})^{⊥} ⊕ (V^{g^{−1}h})^{⊥} = (V^{h})^{⊥} .

This induces a binary relation ≤ on the quotient group G/K as well: For g, h in G, define gK ≤ hK when g ≤ h. Note the relation does not depend on choice of representatives of cosets, as V^{g} = V^{h} whenever gK = hK for g, h in G.

The relation ≤ appears in work of Brady and Watt [4] on orthogonal transformations. Their arguments apply equally well to our setting of isometries with respect to some inner product and the quotient group G/K. (Note that if G does not act faithfully, the binary relation ≤ on G may not be anti-symmetric and thus may not define a partial order on G.)

Lemma 2.3 (Brady and Watt [4]). The relation ≤ is a partial order on G/K. If G acts faithfully, the relation ≤ is a partial order on G.

We shall use the following elements in the sequel.

Definition 2.4. For each g ∈ G, let vol_{g}^{⊥} be a choice of nonzero element in the one-dimensional space ⋀^{codim V^{g}}((V^{g})^{⊥})^{∗}.

We show in the next lemma how these choices determine a multiplicative cocycle. A function ϑ : G × G → ℂ is a multiplicative 2-cocycle on G if

ϑ(gh, k)ϑ(g, h) = ϑ(g, hk)ϑ(h, k)

for all g, h, k ∈ G. We may use any such cocycle to define a new algebra structure on the group algebra ℂ[G], a generalization of a twisted group algebra (in which the values of ϑ may include 0): Let ℂ^{ϑ}[G] be the ℂ-algebra with basis G and multiplication g ·_{ϑ} h = ϑ(g, h) gh for all g, h ∈ G. Associativity is equivalent to the 2-cocycle identity. If ϑ(g, 1_{G}) = 1 = ϑ(1_{G}, g) for all g ∈ G, where 1_{G} denotes the identity element of G, then ℂ^{ϑ}[G] has multiplicative identity 1_{G}.

We canonically embed each space ⋀((V^{g})^{⊥})^{∗} into ⋀ V^{∗}.

Proposition 2.5. For all g and h in G,

vol_{g}^{⊥} ∧ vol_{h}^{⊥} = ϑ(g, h) vol_{gh}^{⊥}

in ⋀ V^{∗} where ϑ : G × G → ℂ is a (multiplicative) 2-cocycle on G with

ϑ(g, h) ≠ 0 if and only if g ≤ gh .

Under wedge product, the algebra ⋀{vol_{g}^{⊥} : g ∈ G} is isomorphic to the (generalized) twisted group algebra ℂ^{ϑ}[G].

Proof. Let g, h be any pair of elements in G. Then V^{g} ∩ V^{h} ⊂ V^{gh}, and hence (V^{gh})^{⊥} ⊂ (V^{g})^{⊥} + (V^{h})^{⊥}. If the sum is direct, then by Lemma 2.1, we have equality of vector subspaces: (V^{gh})^{⊥} = (V^{g})^{⊥} ⊕ (V^{h})^{⊥}. If the sum is not direct, then vol_{g}^{⊥} ∧ vol_{h}^{⊥} = 0. In either case, the product vol_{g}^{⊥} ∧ vol_{h}^{⊥} is a (possibly zero) scalar multiple of vol_{gh}^{⊥}. Hence, there is a scalar ϑ(g, h) ∈ ℂ such that

vol_{g}^{⊥} ∧ vol_{h}^{⊥} = ϑ(g, h) vol_{gh}^{⊥} .

Note that ϑ(g, h) ≠ 0 if and only if codim V^{g} + codim V^{h} = codim V^{gh} (by Lemma 2.1) if and only if g ≤ gh. By associativity of the exterior algebra, the function ϑ : G × G → ℂ is a (multiplicative) 2-cocycle on G. □

We shall also need the following easy lemma, which is a consequence of the fact that (V^{g})^{⊥} = im(1 − g) for all g:

Lemma 2.6. For all g in G:

• f − ^{g}f lies in the ideal I((V^{g})^{⊥}) for all f ∈ S(V), and

• dv ∧ vol_{g}^{⊥} = ^{g}dv ∧ vol_{g}^{⊥} for all dv ∈ ⋀^{}V.

3. Skew group algebra and Hochschild cohomology

In this section, we recall the basic definitions of the skew group algebra and Hochschild cohomology, as well as a fundamental theorem describing the cohomology of the skew group algebra as a space of invariants. We work over the complex numbers ℂ.

Let A denote any ℂ-algebra on which G acts by automorphisms. The skew group algebra (or smash product) A#G is the vector space A ⊗ ℂG with multiplication given by

(a ⊗ g)(b ⊗ h) = a(^{g}b) ⊗ gh

for all a, b ∈ A and g, h ∈ G.

The Hochschild cohomology of a ℂ-algebra A (such as S(V) or S(V)#G), with coefficients in an A-bimodule M, is the graded vector space HH^{}(A, M) = Ext_{A^{e}}(A, M), where A^{e} = A ⊗ A^{op} acts on A by left and right multiplication. We abbreviate HH^{}(A) = HH^{}(A, A).

To construct the cohomology HH^{}(A, M), one applies the functor Hom_{A^{e}}(−, M) to a projective resolution of A as an A^{e}-module, for example, to the bar resolution

(3.1)  ⋯ ⟶^{δ_{3}} A^{⊗4} ⟶^{δ_{2}} A^{⊗3} ⟶^{δ_{1}} A^{e} ⟶^{m} A → 0,

where δ_{p}(a_{0} ⊗ ⋯ ⊗ a_{p+1}) = ∑_{j=0}^{p}(−1)^{j}a_{0} ⊗ ⋯ ⊗ a_{j}a_{j+1} ⊗ ⋯ ⊗ a_{p+1} for all a_{0}, …, a_{p+1} ∈ A, and m is multiplication. For all p ≥ 0, Hom_{A^{e}}(A^{(p+2)}, M) ≅ Hom_{ℂ}(A^{⊗p}, M), and we identify these two vector spaces in what follows.

When M is itself an algebra, the Hochschild cohomology HH^{}(A, M) is a graded associative algebra under the cup product, defined at the cochain level on the bar complex (see [14, §7]): Let f ∈ Hom_{ℂ}(A^{⊗p}, M) and f′ ∈ Hom_{ℂ}(A^{⊗q}, M); then the cup product f ⌣ f′ in Hom_{ℂ}(A^{⊗(p+q)}, M) is given by

(f ⌣ f′)(a_{1} ⊗ ⋯ ⊗ a_{p+q}) = f(a_{1} ⊗ ⋯ ⊗ a_{p})f′(a_{p+1} ⊗ ⋯ ⊗ a_{p+q})

for all a_{1}, …, a_{p+q} ∈ A. We seek to describe the algebra structure under cup product explicitly in the case A = S(V), M = S(V)#G, and in the case A = M = S(V)#G. These two cases are related by a well-known theorem that we state next.

Since |G| is invertible, a result of Ştefan [27, Cor. 3.4] implies in our setting that there is a G-action giving an isomorphism of graded algebras (under the cup product):

Theorem 3.2.

HH^{}(S(V)#G) ≅ HH^{}(S(V), S(V)#G)^{G} .

(Specifically, the action of G on V extends naturally to the bar complex of S(V) and thus induces an action on HH^{}(S(V), S(V)#G), from which we define Ginvariant cohomology in the theorem. In fact, any projective resolution of S(V) compatible with the action of G may be used to define the G-invariant cohomology.) We thus concentrate on describing the cup product on HH^{}(S(V), S(V)#G).

4. Koszul and bar resolutions

One may use either the Koszul or the bar resolution of S(V) to describe the cohomology HH^{}(S(V), S(V)#G) and thus its G-invariant subalgebra, HH^{}(S(V)#G). The Koszul resolution is the following free S(V)^{e}-resolution of S(V):

(4.1) ⋯ ⟶^{d_{3}} S(V)^{e} ⊗ ⋀^{2} V ⟶^{d_{2}} S(V)^{e} ⊗ ⋀^{1} V ⟶^{d_{1}} S(V)^{e} ⟶^{m} S(V) → 0,

where the differential d is given by

(4.2) d_{p}(1⊗1⊗v_{j_{1}}∧⋯∧v_{j_{p}}) = ∑_{i=1}^{p}(−1)^{i+1}(v_{j_{i}}⊗1−1⊗v_{j_{i}})⊗(v_{j_{1}}∧⋯∧v̂_{j_{i}}∧⋯∧v_{j_{p}}),

for all v_{j_{1}}, …, v_{j_{p}} ∈ V. Let Φ be the canonical inclusion (a chain map) of the Koszul resolution (4.1) into the bar resolution (3.1):

\xymatrix{⋯ \ar[r] & S(V)^{⊗4} \ar[r]^{δ_{2}} & S(V)^{⊗3} \ar[r]^{δ_{1}} & S(V)^{e} \ar[r]^{m} & S(V) \ar[r] & 0 \\ ⋯ \ar[r] & S(V)^{e} ⊗ ⋀^{2} V \ar[r]^{d_{2}} \ar[u]_{Φ_{2}} & S(V)^{e} ⊗ ⋀^{1} V \ar[r]^{d_{1}} \ar[u]_{Φ_{1}} & S(V)^{e} \ar[r]^{m} \ar[u]_{=} & S(V) \ar[r] \ar[u]_{=} & 0,}

that is, for all p ≥ 1,

Φ_{p} : S(V)^{e} ⊗ ⋀^{p} V → S(V)^{⊗(p+2)},

(4.3) Φ_{p}(1 ⊗ 1 ⊗ v_{j_{1}} ∧ ⋯ ∧ v_{j_{p}}) = ∑_{π∈Sym_{p}} sgn(π) ⊗ v_{j_{π(1)}} ⊗ ⋯ ⊗ v_{j_{π(p)}} ⊗ 1 ,

for all v_{j_{1}}, …, v_{j_{p}} ∈ V, where Sym_{p} denotes the symmetric group on the set {1, …, p}. Note that Φ is invariant under the action of GL(V), i.e., ^{h}Φ = Φ for all h in GL(V).

One finds cohomology HH^{}(S(V), M) by applying the functor Hom_{S(V)^{e}}(−, M) to either of the above two resolutions and dropping the term Hom_{S(V)^{e}}(S(V), M). We make the customary identifications: For each p, set Hom_{S(V)^{e}}(S(V)^{⊗(p+2)}, M) = Hom_{ℂ}(S(V)^{⊗p}, M) as before, and

(4.4)  Hom_{S(V)^{e}}(S(V)^{e} ⊗ ⋀^{p} V, M) = Hom_{ℂ}(⋀^{p} V, M) = ⋀^{p} V^{∗} ⊗ M .

In the case M = S(V)#G, we write S(V) ⊗ ⋀^{p} V^{∗} ⊗ ℂ[G] for the vector space ⋀^{p} V^{∗} ⊗ M.

We obtain a commutative diagram giving two different cochain complexes describing the cohomology HH^{}(S(V), S(V)#G). In [25], we introduced a “combinatorial converter” map Υ (whose definition is recalled in the next section), which serves as an inverse to the induced map Φ^{∗} and converts between complexes:

(4.5)  \xymatrix{Hom_{ℂ}(S(V)^{⊗p}, S(V)#G) \ar[r]^{δ^{∗}} \ar[d]^{Φ_{p}^{∗}} & Hom_{ℂ}(S(V)^{⊗(p+1)}, S(V)#G) \ar[d]^{Φ_{p+1}^{∗}} \\ S(V) ⊗ ⋀^{p} V^{∗} ⊗ ℂ[G] \ar[r]^{d^{∗}} \ar[u]^{Υ_{p}} & S(V) ⊗ ⋀^{p+1} V^{∗} ⊗ ℂ[G] . \ar[u]^{Υ_{p+1}}}

We use the maps Υ and Φ^{∗} in our analysis of the cup product in later sections.

We describe cohomology explicitly in terms of cocycles and coboundaries. Under the identification (4.4), Hochschild cohomology HH^{}(S(V), S(V)#G) arises from the complex of cochains

(4.6)  C^{}:= ⨁_{g∈G} S(V) ⊗ ⋀^{}V^{∗} ⊗ g .

One may determine the set of cocycles and coboundaries explicitly as the kernel and image of the induced map d^{∗} (under the above identifications (4.4)). We set

(4.7)  Z^{}:= ⨁_{g∈G} S(V) ⊗ ⋀^{−codim V^{g}}(V^{g})^{∗} ⊗ ⋀^{codim V^{g}}((V^{g})^{⊥})^{∗} ⊗ g ,

a subspace of the space of cocycles, and

B^{}:= ⨁_{g∈G} I((V^{g})^{⊥}) ⊗ ⋀^{−codim V^{g}}(V^{g})^{∗} ⊗ ⋀^{codim V^{g}}((V^{g})^{⊥})^{∗} ⊗ g ,

a subspace of the space of coboundaries, where I((V^{g})^{⊥}) is the ideal of S(V) generated by (V^{g})^{⊥}. (We agree that a negative exterior power of a space is defined to be 0.) We regard these subspaces Z^{}and B^{}as subsets of the cochains C^{}after making canonical identifications (we identify W_{1} ⊗ W_{2} with W_{1} ∧ W_{2} for any subspaces W_{1}, W_{2} of V intersecting trivially). We refer to cochains, cocycles, and coboundaries as vector forms “tagged” by the group elements indexing the direct summands above.

The next remark explains that we may view Z^{}as a substitute for the set of cocycles and B^{}as a substitute for the set of coboundaries. We use this alternate description of cohomology (as Z^{}/B^{}) in Section 6 to expose a smash product structure.

Remark 4.8. Note that one may use the isomorphism V/(V^{g})^{⊥} ≅ V^{g} to select a set of a representatives of cohomology classes: Let

H^{}:= ⨁_{g∈G} S(V^{g}) ⊗ ⋀^{−codim V^{g}}(V^{g})^{∗} ⊗ ⋀^{codim V^{g}}((V^{g})^{⊥})^{∗} ⊗ g .

Then H^{}≅ Z^{}/B^{}via the canonical map Proj_{H} : Z^{}/B^{}→ H^{}induced from the compositions, for each g in G,

S(V) ⟶ S(V)/I((V^{g})^{⊥}) ⟶^{∼} S(V^{g}) .

Farinati [13] and Ginzburg and Kaledin [15] showed that

HH^{}(S(V), S(V)#G) ≅ H^{}≅ Z^{}/B^{}.

5. Combinatorial converter map

We recall the definition of the combinatorial converter map Υ (in Diagram 4.5) introduced in [25]. A nonidentity element of GL(V) is a reflection if it fixes a hyperplane in V pointwise. Given any basis v_{1}, …, v_{n} of V, let ∂/∂v_{i} denote the usual partial differential operator with respect to v_{i}. In addition, given a complex number ϵ ≠ 1, we define the ϵ-quantum partial differential operator with respect to v := v_{i} as the scaled Demazure (BBG) operator ∂_{v,ϵ} : S(V) → S(V) given by

(5.1)  ∂_{v,ϵ}(f) = (1 − ϵ)^{−1} \frac{f − ^{s}f}{v} = \frac{f − ^{s}f}{v − ^{s}v} ,

where s ∈ GL(V) is the reflection whose matrix with respect to the basis v_{1}, …, v_{n} is diag(1, …, 1, ϵ, 1, …, 1) with ϵ in the i-th slot. Set ∂_{v,ϵ} = ∂/∂v when ϵ = 1. The operator ∂_{v,ϵ} coincides with the usual definition of quantum partial differentiation: One takes the ordinary partial derivative with respect to v but instead of multiplying each monomial by its degree k in v, one multiplies by the quantum integer

[k]_{ϵ} := 1 + ϵ + ϵ^{2} + ⋯ + ϵ^{k−1}.

We define the map Υ in terms of these Demazure operators. For each g in G, fix a basis B_{g} = {v_{1}, …, v_{n}} of V consisting of eigenvectors of g with corresponding eigenvalues ϵ_{1}, …, ϵ_{n}. Decompose g into (commuting) reflections diagonal in this basis: Let g = s_{1} ⋯ s_{n} where each s_{i} is either the identity or a reflection defined by s_{i}(v_{j}) = v_{j} for j ≠ i and s_{i}(v_{i}) = ϵ_{i}v_{i}. Let ∂_{i} := ∂_{v_{i},ϵ_{i}}, the quantum partial derivative with respect to B_{g}. Recall that C^{}denotes cochains (see (4.6)).

Definition 5.2. We define a map Υ from the dual Koszul complex to the dual bar complex with coefficients in S(V)#G:

Υ_{p} : C^{p} → Hom_{ℂ}(S(V)^{⊗p}, S(V)#G) .

For g in G with basis B_{g} = {v_{1}, …, v_{n}} of V as above, and α = f_{g}⊗v_{j_{1}}^{∗}∧⋯∧v_{j_{p}}^{∗}⊗g with f_{g} ∈ S(V) and 1 ≤ j_{1} < … < j_{p} ≤ n, define Υ(α) : S(V)^{⊗p} → S(V)#G by

Υ(α)(f_{1} ⊗ ⋯ ⊗ f_{p}) = ( ∏_{k=1,...,p} ^{s_{1}s_{2}···s_{j_{k}−1}}(∂_{j_{k}}f_{k}))f_{g} ⊗ g .

By Theorem 5.5 below, Υ is a cochain map. Thus Υ induces a map on the cohomology HH^{}(S(V), S(V)#G), which we denote by Υ as well.

We make the following remark, which will be needed in our analysis of the cup product in Section 7.

Remark 5.3. For the fixed basis B_{g} = {v_{1}, …, v_{n}} and α = f_{g} ⊗ v_{j_{1}}^{∗} ∧ ⋯ ∧ v_{j_{p}}^{∗} ⊗ g in C^{p} (with j_{1} < … < j_{p}), note that

Υ(α)(v_{i_{1}} ⊗ ⋯ ⊗ v_{i_{p}}) = 0 unless i_{1} = j_{1}, …, i_{p} = j_{p} .

In general, Υ(α)(f_{1} ⊗ ⋯ ⊗ f_{p}) = 0 whenever \frac{∂}{∂v_{j_{k}}} (f_{k}) = 0 for some k.

We shall use the following consequence of the definitions.

Proposition 5.4. For any choices of bases defining Υ,

Φ^{∗}Υ = 1

as a map on cochains C^{}.

Proof. Consider a nonzero cochain α in C^{p}. Without loss of generality, suppose that α = f_{g} ⊗ v_{1}^{∗} ∧ ⋯ ∧ v_{p}^{∗} ⊗ g for some g in G, where f_{g} is in S(V) and B_{g} = {v_{1}, …, v_{n}} is the fixed basis of eigenvectors of g. Then for all 1 ≤ i_{1}, …, i_{p} ≤ n,

(Φ^{∗}Υ)(α)(v_{i_{1}} ∧ ⋯ ∧ v_{i_{p}}) = Υ(α)(Φ(v_{i_{1}} ∧ ⋯ ∧ v_{i_{p}}))

= Υ(α) ( ∑_{π∈Sym_{p}} sgn(π) v_{i_{π(1)}} ⊗ ⋯ ⊗ v_{i_{π(p)}})

= ∑_{π∈Sym_{p}} sgn(π) Υ(α)(v_{i_{π(1)}} ⊗ ⋯ ⊗ v_{i_{π(p)}}) ,

while  α(v_{i_{1}} ∧ ⋯ ∧ v_{i_{p}}) = f_{g} (v_{1}^{∗} ∧ ⋯ ∧ v_{p}^{∗})(v_{i_{1}} ∧ ⋯ ∧ v_{i_{p}}) ⊗ g .

By Remark 5.3, both expressions are zero unless {v_{i_{1}}, … v_{i_{p}}} = {v_{1}, …, v_{p}}, in which case both yield sgn(π) f_{g} ⊗g, where π is the permutation sending (i_{1}, …, i_{p}) to (1, …, p). □

We summarize results needed from [25]:

Theorem 5.5. The combinatorial converter map

Υ : Dual Koszul Complex → Dual Bar Complex
C^{p} → Hom_{ℂ}(S(V)^{⊗p}, S(V)#G)

induces isomorphisms of cohomology independent of choices of bases:

• For any basis B_{g} of eigenvectors for any g in G, Υ is a cochain map.

• Although the cochain map Υ depends on choices B_{g} for g in G, the induced map Υ on cohomology HH^{}(S(V), S(V)#G) is independent of choices.

• The map Υ induces an automorphism of HH^{}(S(V), S(V)#G) with inverse automorphism Φ^{∗}. Specifically, Υ and Φ^{∗} convert between expressions of cohomology in terms of the Koszul resolution and the bar resolution.

• The map Υ on HH^{}(S(V), S(V)#G) is G-invariant and hence induces an automorphism on HH^{}(S(V)#G) ≅ HH^{}(S(V), S(V)#G)^{G} .

Remark 5.6. We do not symmetrize Υ, in comparison with similar maps in the literature (see Anno [1] and Halbout and Tang [17]). Since they are chain maps, our maps are the same on cohomology as their symmetrized versions. Symmetrization may be more elegant, however unsymmetrized maps can be more convenient for computation.

6. Smash product structure

In this section, we transform the “cohomology of a group action” into a “group action on cohomology” by viewing both as algebras. We relate Hochschild cohomology of the skew group algebra, HH^{}(S(V)#G), to the skew group algebra of a Hochschild cohomology algebra, HH^{}(S(V))#G: We manifest the first algebra as a subquotient of the second. We thus twist the cup product on HH^{}(S(V)) by the group action and obtain a natural smash product on HH^{}(S(V)#G). In the next section, we show that the cup product on HH^{}(S(V)#G) is precisely this natural smash product.

We first embed HH^{}(S(V), S(V)#G) as a graded vector space into

HH^{}(S(V)) # G ,

the skew group algebra determined by the action of G on HH^{}(S(V)) induced from its action on V. Note that for a general algebra S with action of G by automorphisms, HH^{}(S, S#G) is a G-graded algebra; we show that in the special case S = S(V), it is not only G-graded, but is very close to being a smash product itself. To see this, we first identify the Hochschild cohomology HH^{}(S(V)) with the set of vector forms on V (cf. [31]):

HH^{}(S(V)) = S(V) ⊗ ⋀^{}V^{∗} .

The group G acts on this tensor product diagonally, and the skew group algebra HH^{}(S(V))#G is the ℂ-vector space of cochains,

HH^{}(S(V)) ⊗ ℂ[G] = S(V) ⊗ ⋀^{}V^{∗} ⊗ ℂ[G] = C^{},

together with smash product

(6.1)  (f_{g} ⊗ dv_{g} ⊗ g) ⋄ (f_{h} ⊗ dv_{h} ⊗ h) = f_{g} ^{g}f_{h} ⊗ (dv_{g} ∧ ^{g}dv_{h}) ⊗ gh ,

where f_{g}, f_{h} ∈ S(V), g, h ∈ G, and dv_{g}, dv_{h} ∈ ⋀^{}V^{∗}.

We regard HH^{}(S(V), S(V)#G) as a vector space subquotient of HH^{}(S(V))#G by identifying with Z^{}/B^{}(see Remark 4.8 and the comments before it):

(6.2) HH^{}(S(V), S(V)#G) = Z^{}/B^{}⊂ C^{}/B^{}= (HH^{}(S(V)) ⊗ ℂ[G])/B^{}.

We next recognize this subquotient of vector spaces as a subquotient of algebras under the smash product.

Proposition 6.3. Under the smash product of HH^{}(S(V))#G, the space Z^{}forms a subalgebra of C^{}and the space B^{}forms an ideal of C^{}: For all cochains α and β in HH^{}(S(V))#G,

• If α and β lie in Z^{}, then α ⋄ β also lies in Z^{},

• If α lies in Z^{}and β lies in B^{}, then α ⋄ β and β ⋄ α also lie in B^{}.

Proof. Let α and β be cocycles in Z^{}. Without loss of generality, suppose that

α = f_{g} ⊗ dv_{g} ⊗ vol_{g}^{⊥} ⊗g and β = f_{h} ⊗ dv_{h} ⊗ vol_{h}^{⊥} ⊗h

for g, h in G, f_{g}, f_{h} in S(V), and dv_{g}, dv_{h} in ⋀ V^{∗}. By Lemma 2.6 (see (6.1)),

α ⋄ β = f_{g} ^{g}f_{h} ⊗ (dv_{g} ∧ vol_{g}^{⊥} ∧^{g}(dv_{h} ∧ vol_{h}^{⊥})) ⊗ gh
= f_{g} ^{g}f_{h} ⊗ (dv_{g} ∧ vol_{g}^{⊥} ∧ dv_{h} ∧ vol_{h}^{⊥}) ⊗ gh
= ± f_{g} ^{g}f_{h} ⊗ (dv_{g} ∧ dv_{h} ∧ vol_{g}^{⊥} ∧ vol_{h}^{⊥}) ⊗ gh .

Assume this product is nonzero. Then by Proposition 2.5 and Lemma 2.1,

(V^{g})^{⊥} ⊕ (V^{h})^{⊥} = (V^{gh})^{⊥} = (V^{hg})^{⊥}

and vol_{g}^{⊥} ∧ vol_{h}^{⊥} is a scalar multiple of vol_{gh}^{⊥}. Hence α ⋄ β lies in Z^{}.

Now assume further that β is a coboundary in B^{}, i.e., that f_{h} lies in the ideal I((V^{h})^{⊥}) of S(V). Note that ^{g}((V^{h})^{⊥}) ⊂ ^{g}((V^{hg})^{⊥}) = (V^{ghgg^{−1}})^{⊥} = (V^{gh})^{⊥}. Hence ^{g}f_{h}, and thus the product f_{g} ^{g}f_{h}, lies in I((V^{gh})^{⊥}). Therefore, α ⋄ β is an element of B^{}. The argument for β ⋄ α is similar (and easier). □

The proposition above immediately implies that the smash product on the skew group algebra HH^{}(S(V))#G induces a smash algebra product on the cohomology HH^{}(S(V), S(V)#G), as we see in the next two results.

Corollary 6.4. The vector space subquotient Z^{}/B^{}of HH^{}(S(V))#G is an algebra subquotient (subalgebra of a quotient of algebras) under the induced smash product.

Note that cohomology classes (and cocycles) in general are represented by sums of elements of the form of α and β given in the next proposition.

Proposition 6.5. The cohomology HH^{}(S(V), S(V)#G) identifies naturally as a graded vector space with an algebra subquotient of the smash product HH^{}(S(V))#G. Under this identification, HH^{}(S(V), S(V)#G) inherits the smash product: For cohomology classes in HH^{}(S(V), S(V)#G) represented by cocycles in Z^{},

α = ∑_{g∈G} f_{g} ⊗ dv_{g} ⊗ g and β = ∑_{h∈G} f_{h} ⊗ dv_{h} ⊗ h ,

where each f_{g}, f_{h} ∈ S(V) and each dv_{g}, dv_{h} ∈ ⋀^{}V^{∗}, the smash product

(6.6)  α ⋄ β = ∑_{{g,h∈G \atop g≤gh}} f_{g} ^{g}f_{h} ⊗ (dv_{g} ∧ ^{g}dv_{h}) ⊗ gh

is also a cocycle representing a class of HH^{}(S(V), S(V)#G).

Proof. We saw in (6.2) that HH^{}(S(V), S(V)#G) is isomorphic to the vector space subquotient

Z^{}/B^{}of (HH^{}(S(V))#G)/B^{}.

(In fact, we may identify HH^{}(S(V), S(V)#G) with a subset of HH^{}(S(V))#G; see Remarks 4.8 and 6.7.) By Proposition 6.3, the smash product on HH^{}(S(V))#G may be restricted to the subset Z^{}and induces a multiplication on the subquotient Z^{}/B^{}(Corollary 6.4). Hence HH^{}(S(V), S(V)#G) is isomorphic to an algebra carrying a natural smash product, and thus it inherits (under this isomorphism) a natural smash product of its own. Formula (6.1) gives a cohomology class representative of the smash product of two cocycles. □

The proposition above explains that the vector space inclusion of (6.2) yields an injection of algebras: The algebra HH^{}(S(V), S(V)#G), under smash product ⋄, is isomorphic to an algebra subquotient of HH^{}(S(V))#G. But for any ℂalgebra A carrying an action of G, the smash product on the skew group algebra A#G maps G-invariants to G-invariants. Hence, the algebra HH^{}(S(V)#G) also inherits a smash product and is isomorphic (under smash product) to an algebra subquotient of HH^{}(S(V))#G. We shall see in Corollary 7.6 below that the same is true under cup product.

Remark 6.7. One may identify HH^{}(S(V), S(V)#G) with the subset H^{}of the smash product HH^{}(S(V))#G by fixing a set of cohomology class repesentatives as in Remark 4.8. But note that H^{}is not closed under the smash product as a subset of HH^{}(S(V))#G—we must take a quotient by coboundaries and again take chosen representatives for this quotient. The induced smash product explicitly becomes

(α, β) ↦ Proj_{H}(α ⋄ β) .

7. Equivalence of cup and smash products

In the previous section, we twisted the cup product on HH^{}(S(V)) by the action of G in a natural way to define a multiplication on HH^{}(S(V), S(V)#G) and also on HH^{}(S(V)#G). In fact, we showed that as a graded vector space, the Hochschild cohomology HH^{}(S(V)#G) maps isomorphically to an algebra subquotient of HH^{}(S(V))#G and thus inherits a natural smash product structure. We now regard HH^{}(S(V), S(V)#G) as an algebra under three operations:

• the cup product ⌣ induced from the bar resolution of S(V),

• a “smash” product ⋄ induced from HH^{}(S(V))#G, and

• the usual multiplication in the algebra tensor product HH^{}(S(V)) ⊗ ℂ[G].

We show in this section that these three basic algebraic operations coincide. This allows us to describe generators for Hochschild cohomology in the next section. The equality of cup and smash products (given in the theorem below) also explains how the cup product, defined on the bar resolution, may be expressed in terms of the Koszul resolution.

Theorem 7.1. For all cocycles α and β in Z^{},

α ⌣ β = α ⋄ β .

On the Hochschild cohomology HH^{}(S(V), S(V)#G), this product is induced from the usual multiplication on the tensor product HH^{}(S(V)) ⊗ ℂ[G] of algebras.

Proof. We need only verify the statement for cocycles in Z^{}of the form α = f_{g} ⊗ dv_{g} ⊗ g and β = f_{h} ⊗ dv_{h} ⊗ h, for some g, h in G, where f_{g}, f_{h} ∈ S(V) and dv_{g} ∈ ⋀^{p} V^{∗}, dv_{h} ∈ ⋀^{q} V^{∗} (as any cocycle in Z^{}will be the sum of such elements).

As Φ^{∗} and Υ are inverse maps on the cohomology HH^{}(S(V), S(V)#G) converting between Koszul and bar cochain complexes (see Diagram 4.5),

α ⌣ β = Φ^{∗}(Υ(α) ⌣ Υ(β))

where ⌣ (on the right hand side) denotes cup product on the bar complex. Suppose B_{g} = {v_{1}, …, v_{n}}, a basis of eigenvectors of g (see Definition 5.2). Without loss of generality, assume that dv_{g} = v_{1}^{∗} ∧ ⋯ ∧ v_{p}^{∗} where the span of v_{1}, …, v_{p} includes (V^{g})^{⊥} (see (4.7)). (In general, dv_{g} will be a sum of such elements with indices relabeled.) By the second part of Lemma 2.6,

α ⋄ β = (f_{g} ⊗ dv_{g} ⊗ g) ⋄ (f_{h} ⊗ dv_{h} ⊗ h)
= f_{g} ^{g}f_{h} ⊗ (dv_{g} ∧ ^{g}dv_{h}) ⊗ gh
= f_{g} ^{g}f_{h} ⊗ (dv_{g} ∧ dv_{h}) ⊗ gh .

We compare the values of α ⌣ β and α ⋄ β on any v_{i_{1}} ∧ ⋯ ∧ v_{i_{p+q}}. Now

(α ⋄ β)(v_{i_{1}} ∧ ⋯ ∧ v_{i_{p+q}}) = f_{g} ^{g}f_{h} (dv_{g} ∧ dv_{h})(v_{i_{1}} ∧ ⋯ ∧ v_{i_{p+q}}) ⊗ gh while

(α ⌣ β)(v_{i_{1}} ∧ ⋯ ∧ v_{i_{p+q}})

= Φ^{∗}(Υ(α) ⌣ Υ(β))(v_{i_{1}} ∧ ⋯ ∧ v_{i_{p+q}})

= (Υ(α) ⌣ Υ(β))( ∑_{π∈Sym_{p+q}} sgn(π)v_{i_{π(1)}} ⊗ ⋯ ⊗ v_{i_{π(p+q)}})

= ∑_{π∈Sym_{p+q}} sgn(π) Υ(α)(v_{i_{π(1)}} ⊗ ⋯ ⊗ v_{i_{π(p)}}) Υ(β)(v_{i_{π(p+1)}} ⊗ ⋯ ⊗ v_{i_{π(p+q)}}) .

By Remark 5.3, both α ⌣ β and α⋄β are readily seen to be zero on v_{i_{1}} ∧⋯∧v_{i_{p+q}} unless {v_{i_{1}}, …, v_{i_{p+q}}} contains {v_{1}, …, v_{p}}. Thus, we may assume (after relabeling indices and also possibly changing signs throughout) that v_{1} = v_{i_{1}}, …, v_{p+q} = v_{i_{p+q}}.

Again by Remark 5.3, (α ⌣ β)(v_{1} ∧ ⋯ ∧ v_{p+q}) is equal to

∑_{π∈Sym_{p}} sgn(π) Υ(α)(v_{π(1)} ⊗ ⋯ ⊗ v_{π(p)}) ∑_{π∈Sym_{q}} sgn(π) Υ(β)(v_{p+π(1)} ⊗ ⋯ ⊗ v_{p+π(q)}) .

Proposition 5.4 then implies that this value of the cup product is just

Φ^{∗}(Υ(α))(v_{1} ∧ ⋯ ∧ v_{p}) · Φ^{∗}(Υ(β))(v_{p+1} ∧ ⋯ ∧ v_{p+q})
= α(v_{1} ∧ ⋯ ∧ v_{p}) · β(v_{p+1} ∧ ⋯ ∧ v_{p+q})
= f_{g} ^{g}f_{h} dv_{g}(v_{1} ∧ ⋯ ∧ v_{p}) dv_{h}(v_{p+1} ∧ ⋯ ∧ v_{p+q}) ⊗ gh
= f_{g} ^{g}f_{h} (dv_{g} ∧ dv_{h})(v_{1} ∧ ⋯ ∧ v_{p+q}) ⊗ gh
= (α ⋄ β)(v_{1} ∧ ⋯ ∧ v_{p+q})

as elements of S(V)#G. Thus α ⌣ β and α ⋄ β agree as cochains.

If this product is nonzero, then f_{h} − ^{g}f_{h} lies in the ideal I((V^{g})^{⊥}) ⊂ I((V^{gh})^{⊥}) by Lemma 2.1, Proposition 2.5, and the first part of Lemma 2.6. Hence,

f_{g} ^{g}f_{h} ⊗ (dv_{g} ∧ dv_{h}) ⊗ gh and f_{g} f_{h} ⊗ (dv_{g} ∧ dv_{h}) ⊗ gh

represent the same class in the cohomology HH^{}(S(V), S(V)#G) = Z^{}/B^{}. Thus the usual multiplication in the tensor product of algebras HH^{}(S(V)) ⊗ ℂ[G] gives the cup product on cohomology. □

We next give an example to show that the cup and smash products do not agree on arbitrary cochains.

Example 7.2. Let G = ℤ/2ℤ × ℤ/2ℤ × ℤ/2ℤ, generated by elements a_{1}, a_{2}, a_{3}. Let V = ℂ^{3} with basis v_{1}, v_{2}, v_{3} on which G acts as follows:

a_{i}v_{j} = (−1)^{δ_{ij}}v_{j}.

Let α = 1 ⊗ v_{3}^{∗} ⊗ a_{1}a_{3} and β = 1 ⊗ v_{1}^{∗} ∧ v_{2}^{∗} ⊗ a_{1}a_{2} in C^{}. Then

α ⋄ β = −1 ⊗ v_{1}^{∗} ∧ v_{2}^{∗} ∧ v_{3}^{∗} ⊗ a_{2}a_{3} = −α ⌣ β.

Our results imply the following explicit formula for the cup product, first given by Anno [1], expressed here in terms of the poset in Definition 2.2 and the multiplicative cocycle ϑ : G×G → ℂ of Proposition 2.5. Note that the condition g < gh in the sum below is included merely for computational convenience, as ϑ(g, h) is nonzero if and only if g < gh. Also note that cohomology classes (and cocycles) in general are represented by sums of such α and β given in the theorem below.

Theorem 7.3. Consider cohomology classes in HH^{}(S(V), S(V)#G) represented by

α = ∑_{g∈G} f_{g} ⊗ dv_{g} ⊗ vol_{g}^{⊥} ⊗g and β = ∑_{h∈G} f_{h} ⊗ dv_{h} ⊗ vol_{h}^{⊥} ⊗h

in Z^{}of degrees p and q, resp. (see (4.7)). Set m = (codim V^{g})(q − codim V^{h}). The cup product α ⌣ β is represented by the cocycle

(7.4)  ∑_{{g,h∈G \atop with g≤gh}}(−1)^{m} ϑ(g, h) f_{g} f_{h} ⊗ (dv_{g} ∧ dv_{h}) ⊗ vol_{gh}^{⊥} ⊗ gh .

Proof. The formula follows directly from Proposition 2.5, Proposition 6.5, and Theorem 7.1 (see its proof). Note that the factor (−1)^{m} arises when we replace vol_{g}^{⊥} ∧ dv_{h} with dv_{h} ∧ vol_{g}^{⊥}. □

Example 7.5. Let G, V be as in Example 7.2. Let α = 1 ⊗ v_{1}^{∗} ⊗ a_{1}, β = 1 ⊗ v_{3}^{∗} ∧ v_{2}^{∗} ⊗ a_{2} in H^{}. Let vol_{a_{1}}^{⊥} = v_{1}^{∗}, vol_{a_{2}}^{⊥} = v_{2}^{∗}, and vol_{a_{1}a_{2}}^{⊥} = v_{1}^{∗} ∧ v_{2}^{∗}, so that ϑ(a_{1}, a_{2}) = 1. Then

α ⌣ β = α ⋄ β = (−1)^{1(2−1)} ⊗^{a_{1}}(v_{3}^{∗}) ∧ v_{1}^{∗} ∧ v_{2}^{∗} ⊗ a_{1}a_{2}
= −1 ⊗ v_{3}^{∗} ∧ vol_{a_{1}a_{2}}^{⊥} ⊗ a_{1}a_{2}.

Corollary 7.6. The algebra HH^{}(S(V), S(V)#G) is isomorphic to an algebra subquotient of HH^{}(S(V))#G. The algebra HH^{}(S(V)#G) is isomorphic to its Ginvariant subalgebra. Hence, it is also an algebra subquotient of HH^{}(S(V))#G.

Proof. The space HH^{}(S(V), S(V)#G) may be written as a (graded) subspace of a vector space quotient of HH^{}(S(V))#G (see (6.2)). Proposition 6.5 states that this subquotient is actually a subquotient of algebras under the smash product (see Proposition 6.3 and Corollary 6.4). But the cup product on HH^{}(S(V), S(V)#G) is the same as this induced smash product by Theorem 7.1. By Theorem 3.2, the cohomology algebra HH^{}(S(V)#G) is isomorphic to the G-invariant subalgebra of HH^{}(S(V), S(V)#G) (both spaces regarded as algebras under their respective cup products). Hence the algebra HH^{}(S(V)#G) is also a subquotient of HH^{}(S(V))#G. □

8. Volume algebra

Results in the last section reveal an interesting subalgebra of the Hochschild cohomology HH^{}(S(V), S(V)#G) isomorphic to algebras that have appeared in the literature before (see Remark 8.2 below); we call it the volume algebra. We use this subalgebra to give algebra generators of cohomology in the next section.

We define the volume algebra with the next proposition. Recall that for each g in G, the form vol_{g}^{⊥} is a choice of nonzero element in the one-dimensional space ⋀^{codim V^{g}}((V^{g})^{⊥})^{∗}. We show that the elements vol_{g}^{⊥} ⊗ g := 1 ⊗ vol_{g}^{⊥} ⊗ g in HH^{}(S(V), S(V)#G) generate a subalgebra which captures the binary relation ≤ on G and reflects the poset structure of G/K (see Definition 2.2). (Note that if k acts trivially on V, then vol_{k}^{⊥} = 1 up to a nonzero constant.)

Proposition 8.1. The ℂ-vector space

A_{vol} := Span_{ℂ}{vol_{g}^{⊥} ⊗ g | g ∈ G}

is a subalgebra of HH^{}(S(V), S(V)#G). The induced multiplication on this subalgebra is given by

(vol_{g}^{⊥} ⊗ g)(vol_{h}^{⊥} ⊗ h) = ϑ(g, h) vol_{gh}^{⊥} ⊗ gh

where ϑ : G × G → ℂ is a (multiplicative) cocycle on G. For any g, h in G, the above product is nonzero (ϑ(g, h) ≠ 0) if and only if g ≤ gh. The algebra A_{vol} is isomorphic to the (generalized) twisted group algebra ℂ^{ϑ}[G].

Proof. Lemma 2.6 and Proposition 2.5 imply that for any pair g, h in G,

(vol_{g}^{⊥} ⊗ g) ⋄ (vol_{h}^{⊥} ⊗ h) = (vol_{g}^{⊥} ∧ ^{g}(vol_{h}^{⊥})) ⊗ gh
= (vol_{g}^{⊥} ∧ vol_{h}^{⊥}) ⊗ gh
= ϑ(g, h) vol_{gh}^{⊥} ⊗ gh ,

for a cocycle ϑ which is nonzero on the pair (g, h) exactly when g ≤ gh. Hence by Theorem 7.1, A_{vol} is closed under cup product and isomorphic to ℂ^{ϑ}[G]. □

The above proposition explains how A_{vol} algebraically captures the geometric relation on G given by ≤. Indeed, the space A_{vol} forms a graded algebra where

deg(vol_{g}^{⊥} ⊗ g) = codim V^{g} for each g ∈ G .

Moreover, for all g, h in G,

g ≤ h if and only if vol_{g}^{⊥} ⊗ g divides vol_{h}^{⊥} ⊗ h in A_{vol} .

Let S be the subset {vol_{g}^{⊥} ⊗ g : g ∈ G} of A_{vol}. Then S is a poset under division:

vol_{g}^{⊥} ⊗ g ≤_{S} vol_{h}^{⊥} ⊗ h if and only if vol_{g}^{⊥} ⊗ g divides vol_{h}^{⊥} ⊗ h in A_{vol} .

If G acts faithfully, we have isomorphic posets:

(G, ≤) ≅ (S, ≤_{S}) .

Remark 8.2. The subalgebra A_{vol} of HH^{}(S(V), S(V)#G) appears in other arenas. If V is a symplectic vector space and G ⊂ Sp(V), the algebra A_{vol} is isomorphic to the graded algebra gr_{F} ℂ[G] associated to the filtration F on G assigning to each group element g the codimension of V^{g}. Its G-invariant subalgebra A_{vol}^{G} is isomorphic to the orbifold cohomology of V/G in this case (see Fantechi and Göttsche [12, §2] or Ginzburg-Kaledin [15]) and also to the Hochschild cohomology of the G-invariant subalgebra of the Weyl algebra A(V) (see Suarez-Alvarez [28]):

HH^{}(A(V)^{G}) ≅ HH^{}(A(V)#G) ≅ A_{vol}^{G} ≅ H_{orb}(V/G) .

Work of several authors has shown that for G = Sym_{n} acting symplectically on ℂ^{2n}, this orbifold cohomology is isomorphic to the cohomology of a Hilbert scheme which is a crepant resolution of the orbifold [12, 20, 30]. Lehn and Sorger gave a description of A_{vol}^{Sym_{n}} in terms of generators and relations [20, Remark 6.3].

9. Cohomology generators given by the poset

In this section, we give generators for the algebra HH^{}(S(V), S(V)#G) in terms of the partial order ≤ on the quotient group G/K, where K is the kernel of the representation of the group G acting on V. In Section 7, we showed that the cup product and an induced smash product on the algebra HH^{}(S(V), S(V)#G) agree. Hence, we simply discuss generation of cohomology as an algebra, without explicit reference to the product. We explain in the next three results how generators for HH^{}(S(V), S(V)#G) are tagged by K together with minimal elements in the poset G/K. Actually, 1_{G/K} := K is the unique minimal element in the poset G/K, and we remove it before seeking minimal elements.

Theorem 9.1. The subalgebra A_{vol} of HH^{}(S(V), S(V)#G) is generated as an algebra over ℂ by

{vol_{g}^{⊥} ⊗ g | g ∈ K or g ∈ [G/K] with gK minimal in the poset G/K − {1_{G/K}}} ,

where [G/K] is any set of coset representatives for G/K.

Proof. Assume h ∉ K. We first observe that we may write vol_{h}^{⊥} ⊗ h as the product of two volume forms, one tagged by any other coset representive of hK and the other tagged by an element of K: Suppose hK = h′K for some h′ in G. Then V^{h} = V^{h′} and vol_{h}^{⊥} = vol_{h′}^{⊥} up to a nonzero scalar. Let k = h^{−1}h′, so vol_{k}^{⊥} is a nonzero scalar itself. Then (see Equation 6.1),

(vol_{h}^{⊥} ⊗ h) ⋄ (vol_{k}^{⊥} ⊗ k) = (vol_{h}^{⊥} ∧ ^{h}(vol_{k}^{⊥})) ⊗ hk,

which is a nonzero scalar multiple of vol_{h}^{⊥} ⊗ h′, and thus of vol_{h′}^{⊥} ⊗ h′. We may therefore assume that h itself lies in [G/K].

Suppose that hK is not minimal in the partial order on G/K − {1_{G/K}}. Then there exists g ∈ [G/K] with K ≠ gK ≠ hK and gK < hK, i.e., (V^{g})^{⊥}⊕(V^{g^{−1}h})^{⊥} = (V^{h})^{⊥}. By Theorem 7.3,

(vol_{g}^{⊥} ⊗ g) ⋄ (vol_{g^{−1}h}^{⊥} ⊗ g^{−1}h) = ϑ(g, g^{−1}h) vol_{h}^{⊥} ⊗ h ,

which is nonzero since gK < hK. By symmetry in the definition of the partial order, g^{−1}hK < hK in the poset G/K − {1_{G/K}} as well. Hence, we have written vol_{h}^{⊥} ⊗h as a product of volume forms each tagged by a group element less than h in the partial order. As the set G/K − {1_{G/K}} is finite, the partial order is well-founded, i.e., every descending chain contains a minimal element. Hence, by induction, vol_{h}^{⊥} ⊗h may be written (up to a scalar) as the product of elements in the set given in the statement of the theorem. □

We now turn to the task of describing generators for the full Hochschild cohomology HH^{}(S(V), S(V)#G) as an algebra. We regard HH^{}(S(V)) as a subalgebra of HH^{}(S(V), S(V)#G) by identifying

HH^{}(S(V)) = S(V) ⊗ ⋀^{}V^{∗} with S(V) ⊗ ⋀^{}V^{∗} ⊗ 1_{G} .

Then HH^{}(S(V), S(V)#G) becomes a module over HH^{}(S(V)) under cup product.

Theorem 9.2. The Hochschild cohomology algebra HH^{}(S(V), S(V)#G) is generated by its subalgebras HH^{}(S(V)) and A_{vol}.

Proof. Let α be an arbitrary element of HH^{}(S(V), S(V)#G). Without loss of generality, assume α = f_{h} ⊗ dy ∧ vol_{h}^{⊥} ⊗ h for some h in G and dy ∈ ⋀^{}(V^{h})^{∗}. By Theorem 7.1 (see (7.4)),

α = (f_{h} ⊗ dy ⊗ 1_{G}) ⋄ (vol_{h}^{⊥} ⊗ h)

where f_{h} ⊗ dy ⊗ 1_{G} identifies with f_{h} ⊗ dy in HH^{}(S(V)). Hence, α lies in the product HH^{}(S(V)) · A_{vol}. □

Recall that for k in K, vol_{k}^{⊥} = 1 up to a nonzero constant in ℂ. The last two theorems then imply:

Corollary 9.3. The Hochschild cohomology algebra HH^{}(S(V), S(V)#G) is generated by its subalgebra HH^{}(S(V)) and

{vol_{g}^{⊥} ⊗ g | g ∈ K or g ∈ [G/K] with gK minimal in the poset G/K − {1_{G/K}}} ,

where [G/K] is any set of coset representatives of G/K.

If G acts faithfully on V, then we may simply take minimal elements in the poset G − {1_{G}} (see Lemma 2.3) to obtain a generating set under cup product:

Corollary 9.4. Assume G acts faithfully on V. The Hochschild cohomology algebra HH^{}(S(V), S(V)#G) is generated by HH^{}(S(V)) and

{vol_{g}^{⊥} ⊗ g | g is minimal in the poset G − {1_{G}}}.

10. Reflection groups

In the previous sections, we described the algebraic structure of the Hochschild cohomology HH^{}(S(V), S(V)#G). These results have a special interpretation for reflection groups and Coxeter groups in particular. We are interested in comparing the codimension of the fixed point space V^{g} of a group element g with its “reflection length” in the group.

Recall that a nonidentity element of GL(V) is a reflection if it fixes a hyperplane in V pointwise. A reflection group is a finite group generated by reflections. A reflection group is called a Coxeter group when it is generated by reflections acting on a real vector space. In this section, we restrict ourselves to the case when G is a reflection group. We define a length function with respect to the set of all reflections inside G. (Note that this definition may differ from the length function defined in terms of a fixed choice of generators for the group G, for example, a choice of simple reflections for a Weyl group.)

Definition 10.1. For each g in G, let l(g) be the minimal number k such that g = s_{1} ⋯ s_{k} for some reflections s_{i} in G. We set l(1_{G}) = 0. We call l : G → ℕ the reflection length function (or “absolute length function”) of G.

The reflection length function induces a partial order ≤_{l} on G:

g ≤_{l} h when l(g) + l(g^{−1}h) = l(h) .

The poset formed by the reflection length function plays an important role in the emerging theory of Artin groups of finite type. This theory relies on a key result for Coxeter groups asserting that the closed interval from the identity group element to a Coxeter element forms a lattice. Brady and Watt [5] gave a case-free proof of this fact by relating the two partial orders ≤ and ≤_{l}. The poset ≤_{l} defined from the reflection length function has received attention not only for Coxeter groups, but for other complex reflection groups G as well. One may define a Coxeter element and again consider the interval (often a lattice) from the identity to the Coxeter element, the so-called poset of noncrossing partitions for G. See, for example, Bessis and Reiner [2].

Note that the length of a linear transformation with respect to the ambient group GL(V) coincides with the codimension of the fixed point space: Any element g in the unitary group U(V) can be written as the product of m reflections in GL(V) and no fewer if and only if codim V^{g} = m (see Brady and Watt [4]).

Reflection length in the group G is bounded below by the codimension of the fixed point space:

(10.2)  l(g) ≥ codim V^{g} for all g in G .

Indeed, if l(g) = m and g = s_{1} ⋯ s_{m} is a product of reflections in G, then

V^{g} ⊃ V^{s_{1}} ∩ ⋯ ∩ V^{s_{m}} ;

but each V^{s_{i}} is a hyperplane, so

codim V^{g} ≤ codim(V^{s_{1}} ∩ ⋯ ∩ V^{s_{m}}) ≤ m .

For certain reflection groups, reflection length coincides with codimension of fixed point space. The arguments of Carter [8] for Weyl groups hold for Coxeter groups as well:

Lemma 10.3. Let G be a (finite) Coxeter group. Then reflection length coincides with codimension of fixed point space: For all g in G,

codim V^{g} = l(g) .

The lemma above implies that for Coxeter groups (which act faithfully), the partial order ≤ above (see Definition 2.2) describing the ring structure of Hochschild cohomology coincides with the partial order ≤_{l} induced from the reflection length function. When these two posets agree (i.e., when ≤ = ≤_{l}), we may express the ring structure of Hochschild cohomology in an elegant way. In particular, the Hochschild cohomology algebra HH^{}(S(V), S(V)#G) is generated in degrees 0 and 1, as we see in the next corollary. Note the analogy with the Hochschild-Kostant-Rosenberg Theorem, which implies that the Hochschild cohomology of a smooth commutative algebra is also generated in degrees 0 and 1.

Theorem 10.4. Suppose G is a (finite) reflection group for which the reflection length function gives codimension of fixed point spaces:

l(g) = codim V^{g} for all g in G .

Then HH^{}(S(V), S(V)#G) is generated (as an algebra) in degrees 0 and 1.

Proof. As G is a reflection group, it acts faithfully on V by definition. Let h ≠ 1_{G} in G. Suppose l(h) = codim V^{h} = m > 0. Then h = s_{1} ⋯ s_{m} for some reflections s_{i} in G. Let s = s_{1} and h′ = s_{2} ⋯ s_{m}, so that h′ = s^{−1}h. Since h can not be expressed as a product of fewer than m reflections, codim V^{h′} = l(h′) = m − 1. It follows that

codim V^{s} + codim V^{s^{−1}h} = codim V^{s} + codim V^{h′} = codim V^{h},

and s ≤ h. In fact, s is minimal in the poset G − {1_{G}}: If g ≤ s, then either codim V^{g} = 0 (i.e., g = 1_{G}) or codim V^{g^{−1}s} = 0 (i.e., g = s), as codim V^{s} = 1.

We have shown that the elements of G−{1_{G}} minimal in the partial order ≤ are the reflections. By Corollary 9.4, HH^{}(S(V), S(V)#G) is generated by HH^{}(S(V)) and by all vol_{s}^{⊥} ⊗ s where s is a reflection in G. The elements vol_{s}^{⊥} ⊗ s each have cohomological degree 1, and HH^{}(S(V)) = S(V)⊗⋀^{}V^{∗} is generated as an algebra (under cup product) by HH^{0}(S(V)) ≅ S(V) and HH^{1}(S(V)) ≅ S(V) ⊗ V^{∗}. The statement follows. □

The above corollary applies not only to Coxeter groups, but to other complex reflection groups as well. Let G(r, 1, n) be the infinite family of complex reflection groups, each of which is the symmetry group of a regular (“Platonic”) polytope in complex space V = ℂ^{n}. The group G(r, 1, n) consists of all those n by n complex matrices which have in each row and column a single nonzero entry, necessarily a primitive r-th root of unity. This group is a natural wreath product of the symmetric group and the cyclic group of order r: G ≅ ℤ/rℤ ≀ Sym_{n}. In fact, G(1, 1, n) is the symmetric group Sym_{n} and G(r, 1, n) is the Weyl group of type B_{n}.

Lemma 10.5. For the infinite family G(r, 1, n), the reflection length function coincides with codimension of fixed point spaces:

l(g) = codim V^{g} for all g in G .

Proof. Let ξ be a primitive r-th root of unity in ℂ. Every element in G(r, 1, n) is conjugate to a product h = c_{1} ⋯ c_{k} of disjoint cycles c_{k} of the form

c_{k} = ξ_{j}^{a} (i, i + 1, …, j)

(i.e., h is block diagonal, with k-th block c_{k}) where ξ_{j} := diag(1, …, 1, ξ, 1, …, 1) is the diagonal reflection with ξ in the j-th entry, a ≥ 0, and where (i, i + 1, …, j) is the matrix (in the natural reflection representation) of the corresponding cycle in Sym_{n}. (See, for example, Section 2B of Ram and the first author [23].) Consider a fixed cycle c = c_{k} as above. We may write the cycle (i, i + 1, …, j) as a product of j − i transpositions (reflections) in Sym_{n}, a subgroup of G(r, 1, n). Hence if ξ^{a} = 1, then c may be expressed as the product of j − i reflections in G(r, 1, n) while codim V^{c} = j−i. If ξ^{a} ≠ 1, then c may be expressed as the product of j−i+1 reflections in G(r, 1, n) while codim V^{c} = j −i+1. In either case, l(c) ≤ codim V^{c}. But reflection length is bounded below by codimension, codim V^{c} ≤ l(c), and hence l(c) = codim V^{c}. Then

l(h) ≤ l(c_{1}) + … + l(c_{k}) = codim V^{c_{1}} + … + codim V^{c_{k}} = codim V^{h} ≤ l(h)

and thus l(h) = codim V^{h}. □

As Theorem 10.4 applies to Coxeter groups by Lemma 10.3 and to the infinite family G(r, 1, n) by Lemma 10.5, we have the following analog of the Hochschild-Kostant-Rosenberg Theorem:

Corollary 10.6. Let G be a Coxeter group or the infinite family G(r, 1, n). Then HH^{}(S(V), S(V)#G) is generated as an algebra in degrees 0 and 1.

Note that the two partial orders ≤ and ≤_{l} do not always agree, i.e., that for a complex reflection group G in general, the reflection length function may not give codimension of fixed point spaces:

Example 10.7. Let G be the complex reflection group G(4, 2, 2), the subgroup of G(4, 1, 2) consisting of those matrices with determinant ±1. Let g be the diagonal matrix diag(i, i) where i = \sqrt{−1} with determinant −1. Every reflection in G has determinant −1, and hence g can only be written as the product of an odd number of reflections. Then codim V^{g} = 2 and yet g can not be written as the product of two reflections.

11. Cup product on the invariant subalgebra

In the above sections, we investigated the cup product on the Hochschild cohomology HH^{}(S(V), S(V)#G). In this section, we describe the cup product on HH^{}(S(V)#G), its G-invariant subalgebra, using standard techniques from group theory. Note that generators for the algebra HH^{}(S(V), S(V)#G) may not be invariant under G and hence do not generally yield generators for HH^{}(S(V)#G).

The cup product on HH^{}(S(V)#G) in cohomological degrees 0 and 1 is easy to describe, and follows from an observation of Farinati [13] (see also [24]):

Lemma 11.1. The only group elements in G that contribute to the Hochschild cohomology HH^{}(S(V)#G) are those which act on V with determinant 1.

Note that if G embeds in SL(V), then every g-component is nonzero (as any element in the one-dimensional subspace 1 ⊗ ⋀^{dim V} V^{∗} ⊗ g of C^{}is automatically invariant under the centralizer of g; see (11.3) below).

Lemma 11.1 together with Remark 4.8 immediately implies that the cup product on HH^{}(S(V)#G) in cohomological degrees 0 and 1 is simply the exterior (wedge) product of forms when G acts faithfully (since reflections do not have determinant 1):

Proposition 11.2. Assume G acts faithfully on V, that is, G embeds in GL(V). Then the cup product of elements in HH^{0}(S(V)#G) and HH^{1}(S(V)#G) is given by the cup product on the G-invariant subspaces of HH^{0}(S(V)) and HH^{1}(S(V)):

HH^{0}(S(V)#G) = S(V)^{G} and HH^{1}(S(V)#G) = (S(V) ⊗ V^{∗})^{G} .

When G acts nonfaithfully, we similarly find that only the kernel K of G acting on V contributes to the cohomology HH^{}(S(V)#G) in degrees 0 and 1:

HH^{0}(S(V)#G) = ( ⨁_{k∈K} S(V)⊗k)^{G} and HH^{1}(S(V)#G) = ( ⨁_{k∈K} S(V)⊗V^{∗}⊗k)^{G} .

The cup product in higher degrees is not as transparent. We give a formula in terms of a fixed set C of representatives of the conjugacy classes of G. We extend the isomorphism of Theorem 3.2, HH^{}(S(V)#G) ≅ HH^{}(S(V), S(V)#G)^{G}:

(11.3)  HH^{}(S(V)#G) ≅ ⨁_{g∈C} HH^{}(S(V), S(V) ⊗ g)^{Z(g)},

where Z(g) denotes the centralizer in G of g. The term HH^{}(S(V), S(V) ⊗ g) is isomorphic to the g-component of H^{}of Remark 4.8. The isomorphism identifies an element α of HH^{}(S(V), S(V) ⊗ g)^{Z(g)} with the sum

∑_{h∈[G/Z(g)]}^{h}α,

where [G/Z(g)] denotes a set of representatives of left cosets of Z(g) in G. In the next proposition, we give a formula for the cup product of HH^{}(S(V)#G) expressed in terms of the additive decomposition (11.3).

If A is any algebra with an action of G by automorphisms, and J < L are subgroups of G, we define the transfer map T_{J}^{L} : A^{J} → A^{L} by

T_{J}^{L}(a) = ∑_{h∈[L/J]}^{h}a,

where [L/J] is a set of representatives of the cosets L/J. To prove the following theorem, we use the theory of Green functors applied to this setting of a group action on an algebra.

Theorem 11.4. The cup product on HH^{}(S(V)#G) induces the following product ˙⌣ on ⨁_{g∈C} HH^{}(S(V), S(V) ⊗ g)^{Z(g)} under the isomorphism (11.3): For α ∈ HH^{}(S(V), S(V) ⊗ g)^{Z(g)} and β ∈ HH^{}(S(V), S(V) ⊗ h)^{Z(h)},

α ˙⌣ β = ∑_{x∈D} T_{^{y}Z(g) ∩^{yx}Z(h)}^{Z(k)}(^{y}α ⌣ ^{yx}β),

where D is a set of representatives of the double cosets Z(g)\G/Z(h), and k = k(x) in C and y = y(x) are chosen so that ^{y}g · ^{yx}h = k. The product ^{y}α ⌣ ^{yx}β is the cup product in HH^{}(S(V), S(V)#G)^{^{y}Z(g)∩^{yx}Z(h)}, to which we apply the transfer map T_{^{y}Z(g) ∩^{yx}Z(h)}^{Z(k)} to obtain an element in the k-component HH^{}(S(V), S(V) ⊗ k)^{Z(k)}.

Proof. Let A = HH^{}(S(V), S(V)#G) (considered as a G-algebra). We obtain a standard Green functor by sending a subgroup L of G to the invariant subring A^{L}. The restriction maps (of the functor) are the inclusions, and the transfer maps are as defined above. The component HH^{}(S(V), S(V) ⊗ g)^{Z(g)} is contained in A^{Z(g)}. In the decomposition (11.3), α ∈ HH^{}(S(V), S(V) ⊗ g)^{Z(g)} on the right side is identified with T_{Z(g)}^{G}(α) on the left side, and similarly for β. The formula in the proposition is the standard one for the product of T_{Z(g)}^{G}(α) and T_{Z(h)}^{G}(β) given by the Mackey formula (e.g., see [29, Prop. 1.10], due to Green). □

The product formula in Theorem 11.4 is more than a theoretical observation, but useful in computations, as the next example shows:

Example 11.5. Let V = ℂ^{6} and G = Sym_{3}. Let v_{1}, w_{1}, v_{2}, w_{2}, v_{3}, w_{3} be the standard orthonormal basis of ℂ^{6}, and let G act on V via the permutation representation on {v_{1}, v_{2}, v_{3}} and on {w_{1}, w_{2}, w_{3}}. Choose 1, (12), (123) as representatives of the conjugacy classes of Sym_{3}. By (11.3), HH^{}(S(V)# Sym_{3}) is isomorphic to

HH^{}(S(V))^{Sym^{3}} ⊕ HH^{}(S(V), S(V) ⊗ (12))^{⟨(12)⟩} ⊕ HH^{}(S(V), S(V) ⊗ (123))^{⟨(123)⟩}. These summands are as follows, considering that the actions of (12) and of (123) on the latter two summands, respectively, are trivial:

HH (S(V))^{Sym^{3}} = (S(V) ⊗ ⋀ V^{∗})^{Sym^{3}}
HH (S(V), S(V) ⊗ (12))^{⟨(12)⟩} = S(V^{(12)}) ⊗ ⋀^{−2}(V^{(12)})^{∗} ⊗ (v_{1}^{∗} − v_{2}^{∗}) ∧ (w_{1}^{∗} − w_{2}^{∗}) ⊗ (12)
HH (S(V), S(V) ⊗ (123))^{⟨(123)⟩} = S(V^{(123)}) ⊗ ⋀^{−4}(V^{(123)})^{∗} ⊗

(v_{1}^{∗} − v_{2}^{∗}) ∧ (w_{1}^{∗} − w_{2}^{∗}) ∧ (v_{2}^{∗} − v_{3}^{∗}) ∧ (w_{2}^{∗} − w_{3}^{∗}) ⊗ (123)

where the fixed point spaces are V^{(12)} = Span_{ℂ}{v_{1} + v_{2}, w_{1} + w_{2}, v_{3}, w_{3}} and V^{(123)} = Span_{ℂ}{v_{1} + v_{2} + v_{3}, w_{1} + w_{2} + w_{3}}. The product of an element of HH^{}(S(V))^{Sym^{3}} with an element in any of the three components is given simply by the componentwise product on the exterior algebras and the symmetric algebras. The product of any element of the (123)-component with any element of the (12)- component is 0 as the exterior product is of linearly dependent elements. The product of any two elements of the (123)-component is 0 by degree considerations.

It remains to determine products of pairs of elements from the (12)-component. In the notation of Theorem 11.4, g = (12) = h, and we may take the set D of representatives of the double cosets ⟨(12)⟩\ Sym_{3} /⟨(12)⟩ to be {1, (123)}. If x = 1, we have (12) · (12) = 1, so we take y = 1; in any corresponding product α ⌣ β, the exterior part is a product of linearly dependent elements. Thus the product corresponding to this choice of x must be 0. Now consider the case when x = (123). We have (12) ·^{(123)}(12) = (12)(123)(12)(132) = (123), and so we may take y = 1.

We now have, for any pair α, β ∈ HH^{}(S(V), S(V) ⊗ (12)), the corresponding cup product in HH^{}(S(V)# Sym_{3}) given in terms of the decomposition (11.3):

α ˙⌣ β = T_{{1}}^{⟨ (123) ⟩}(α ⌣ ^{(123)}β).

For example,

vol_{(12)}^{⊥} ˙⌣ vol_{(12)}^{⊥} = T_{{1}}^{⟨ (123) ⟩}(vol_{(12)}^{⊥} ⌣ vol_{(23)}^{⊥}) = 3 vol_{(123)}^{⊥} .

The efficient formula of Theorem 11.4 saves time by reducing nine product calculations to one in this case.

Acknowledgements. The second author thanks Ragnar-Olaf Buchweitz for many helpful conversations, particularly regarding an early version of Theorem 7.3. The authors thank Victor Reiner for aid with Lemma 10.5 and for finding Example 10.7 for them.

References

[1] R. Anno, “Multiplicative structure on the Hochschild cohomology of crossed product algebras,” arxiv:math.QA/0511396.

[2] D. Bessis and V. Reiner, “Cyclic sieving of noncrossing partitions for complex reflection groups,” to appear in Annals of Combinatorics, arxiv:math.QA/0701792.

[3] R. V. Bezrukavnikov and Kaledin, D. B., “McKay equivalence for symplectic resolutions of quotient singularities,” Tr. Mat. Inst. Steklova 246 (2004), Algebr. Geom. Metody, Svyazi i Prilozh., 20–42; translation in Proc. Steklov Inst. Math. 2004, no. 3 (246), 13–33.

[4] T. Brady and C. Watt, “A partial order on the orthogonal group”, Comm. Algebra 30 (2002), no. 8, 3749–3754.

[5] T. Brady and C. Watt, “Non-crossing partition lattices in finite real reflection groups,” Trans. Amer. Math. Soc. 360 (2008), no. 4, 1983–2005.

[6] T. Bridgeland and A. King and M. Reid, “The McKay correspondence as an equivalence of derived categories,” J. Amer. Math. Soc. 14 (2001), no. 3, 535–554.

[7] A. Căldăraru, A. Giaquinto, and S. Witherspoon, “Algebraic deformations arising from orbifolds with discrete torsion,” J. Pure Appl. Algebra 187 (2004), no. 1–3, 51–70.

[8] R. W. Carter, “Conjugacy classes in the Weyl group,” Compositio Math. 25 (1972), 1–59.

[9] T. Chmutova, “Twisted symplectic reflection algebras,” math.RT/0505653.

[10] C. Dezélée, “Une généralisation de l’algèbre de Hecke graduée de type B,” math.RT/0304484.

[11] P. Etingof and V. Ginzburg, “Symplectic reflection algebras, Calogero-Moser space, and deformed Harish-Chandra homomorphism,” Invent. Math. 147 (2002), no. 2, 243–348.

[12] B. Fantechi and L. Göttsche, “Orbifold cohomology for global quotients,” Duke Math. J. 117 (2003), no. 2, 197–227.

[13] M. Farinati, “Hochschild duality, localization, and smash products,” J. Algebra 284 (2005), no. 1, 415–434.

[14] M. Gerstenhaber, “The cohomology structure of an associative ring,” Ann. Math. 78 (1963), no. 2, 267–288.

[15] V. Ginzburg and D. Kaledin, “Poisson deformations of symplectic quotient singularities,” Adv. Math. 186 (2004), no. 1, 1–57.

[16] I. Gordon and S. P. Smith, “Representations of symplectic reflection algebras and resolutions of deformations of symplectic quotient singularities,” Math. Ann. 330 (2004), no. 1, 185–200.

[17] G. Halbout and X. Tang, “Noncommutative Poisson structures on orbifolds,” math.QA/0606436.

[18] J. Harris, Algebraic Geometry: A First Course, Graduate Texts in Mathematics, v. 133, Springer-Verlag, New York, 1995.

[19] B. Keller, “Hochschild cohomology and derived Picard groups,” J. Pure Appl. Algebra 190 (2004), no. 1-3, 177–196.

[20] M. Lehn and C. Sorger, “Symmetric groups and the cup product on the cohomology of Hilbert schemes,” Duke Math. J. 110 (2001), 345–357.

[21] G. Lusztig, “Affine Hecke algebras and their graded version,” J. Amer. Math. Soc. 2 (1989), no. 3, 599–635.

[22] M. J. Pflaum, H. B. Posthuma, X. Tang, and H. H. Tseng, “Orbifold cup products and ring structures on Hochschild cohomologies,” arxiv:math/0706.0027.

[23] A. Ram and A. V. Shepler, “Classification of graded Hecke algebras for complex reflection groups,” Comment. Math. Helv. 78 (2003), 308–334.

[24] A. V. Shepler and S. Witherspoon, “Hochschild cohomology and graded Hecke algebras,” Trans. Amer. Math. Soc. 360 (2008), no. 8, 3975–4005.

[25] A. V. Shepler and S. Witherspoon, “Quantum differentiation and chain maps of bimodule complexes”, submitted.

[26] N. Snashall and Ø. Solberg, “Support varieties and Hochschild cohomology rings,” Proc. London Math. Soc. (3) 88 (2004), no. 3, 705–732.

[27] D. Ştefan, “Hochschild cohomology on Hopf Galois extensions,” J. Pure Appl. Algebra 103 (1995), 221–233.

[28] M. Suarez-Alvarez, “Algebra structure on the Hochschild cohomology of the ring of invariants of a Weyl algebra under a finite group,” J. Algebra 248 (2002), 291–306.

[29] J. Thévenaz, “Some remarks on G-functors and the Brauer morphism,” J. reine angew. Math. 384 (1988), 24–56.

[30] B. Uribe, “Orbifold cohomology of the symmetric product,” Comm. Anal. Geom. 13 (2005), no. 1, 113–128.

[31] C. A. Weibel, An Introduction to Homological Algebra, Cambridge Studies in Adv. Math. 38, Cambridge Univ. Press, Cambridge, 1994.

Department of Mathematics, University of North Texas, Denton, Texas 76203, USA

E-mail address: ashepler@unt.edu

Department of Mathematics, Texas A&M University, College Station, Texas 77843, USA

E-mail address: sjw@math.tamu.edu